# The Dial-a-Ride Problem with Limited Pickups per Trip

### v2

@@ Appendix C Full computational results for MDARP-LPT
$S_{RF}$ $S_{FF}$ $S_{PFF}$ F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap F Net MIP Time OBJ LB Gap a3-18 37 7.5 0.7 8.2 229.2 229.2 0.00% 33 15.1 1.3 16.5 229.2 229.2 0.00% 29 8.0 0.3 8.4 229.2 229.2 0.00% a3-30 36 6.9 0.6 7.6 325.8 325.8 0.00% 36 7.5 0.6 8.2 325.8 325.8 0.00% 34 7.0 2.0 9.2 325.8 325.8 0.00% a3-36 47 9.1 5.9 15.1 397.2 397.2 0.00% 44 6.3 6.5 12.9 397.2 397.2 0.00% 34 15.2 4.0 19.3 397.2 397.2 0.00% a4-48 116 8.3 40.0 48.6 450.8 450.8 0.01% 109 10.0 27.5 37.9 450.8 450.8 0.00% 92 21.1 23.7 45.1 450.8 450.8 0.00% a6-60 175 9.3 766.0 775.9 556.1 556.1 0.00% 155 7.5 490.1 498.0 556.1 556.1 0.00% 113 40.3 396.6 437.4 556.1 556.1 0.00% a7-70 507 12.0 1786.8 1800.0 552.5 549.8 0.48% 482 14.2 861.7 876.9 552.5 552.5 0.00% 250 139.2 504.8 645.1 552.5 552.5 0.00% a8-80 495 16.9 1781.1 1800.0 663.9 641.2 3.41% 586 15.9 1783.0 1800.0 670.7 642.9 4.15% 318 137.9 1661.4 1800.0 661.8 647.6 2.14% Aver-A 201.9 10.0 636.5 0.56% 206.4 10.9 464.3 0.59% 124.3 52.7 423.5 0.31% b3-36 146 8.3 72.9 81.9 389.3 389.3 0.00% 148 9.5 36.9 46.9 389.3 389.3 0.00% 100 15.2 27.6 42.9 389.3 389.3 0.00% b4-40 370 8.7 45.6 54.7 355.7 355.7 0.00% 314 8.6 23.5 32.3 355.7 355.7 0.00% 178 21.0 16.3 37.5 355.7 355.7 0.00% b5-50 1649 16.8 1781.2 1800.0 436.3 427.0 2.12% 1635 20.4 659.2 680.9 436.3 436.3 0.00% 753 45.4 298.9 344.9 436.3 436.3 0.00% b6-60 2269 31.3 1763.9 1800.0 527.9 512.3 2.95% 2041 28.8 1768.7 1800.0 527.7 522.6 0.98% 816 139.1 397.4 537.9 527.7 527.7 0.00% b7-70 1541 18.6 1778.4 1800.0 530.5 509.6 3.93% 2000 32.5 1764.2 1800.0 528.4 519.1 1.76% 747 233.9 1565.0 1800.0 528.4 518.8 1.81% Aver-B 1195.0 16.7 1107.3 1.80% 1227.6 20.0 872.0 0.55% 518.8 90.9 552.6 0.36% 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; MIP: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver. The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively.
@@
Table 2: Computational performance of FFF with different fragment sets for MDARP-LPT within 30 minutes (basic instances: $|V|=\lceil n/3\rceil$, $L=4$)

### v3

@@ Appendix B Full computational results of PSFF for DARP-LPT and MDARP-LPT
$S_{RF}$ $S_{FF}$ $S_{MF}$ $S_{EFF}$ $S_{ERF}$ Name OBJ F Net CPU Time F Net CPU Time F Net CPU Time F Net CPU Time F Net CPU Time 30N_4K_A 409.8 376 0.2 0.1 0.4 372 0.4 0.1 0.5 410 0.4 0.1 0.5 3499 2.5 0.1 2.9 12876 3.7 0.5 5.8 30N_4K_B 482.4 179 0.1 0.1 0.3 171 0.1 0.1 0.4 180 0.2 0.2 0.4 1771 1.5 0.1 1.7 6404 2.2 0.3 3.0 30N_4K_C 436.5 603 0.4 1.2 1.9 899 1.0 1.1 2.2 764 1.2 2.1 3.4 5250 4.8 0.9 6.1 17664 5.1 3.1 9.6 30N_5K_A 425.9 525 0.3 0.5 1.0 759 0.7 0.4 1.4 688 0.9 0.5 1.6 4823 3.5 0.5 4.3 13595 4.0 0.8 6.1 30N_5K_B 450.9 327 0.2 0.5 0.8 353 0.3 0.5 0.9 365 0.5 0.5 1.1 2912 2.1 0.3 2.6 8955 2.6 0.9 4.4 30N_5K_C 460.0 466 0.3 0.7 1.1 547 0.5 0.4 1.0 547 0.7 0.9 1.7 3251 2.9 0.4 3.6 11851 4.0 1.5 6.5 40N_4K_A 554.0 759 0.4 1.0 1.7 965 1.4 0.5 2.1 887 1.4 0.7 2.3 7629 10.3 1.5 12.7 27638 8.1 3.0 13.4 40N_4K_B 552.7 998 0.9 2.0 3.3 1375 1.9 1.3 3.5 1273 1.9 1.5 3.6 14541 13.3 3.8 18.0 52456 15.5 5.3 25.6 40N_4K_C 591.6 629 0.3 0.7 1.2 657 1.3 0.7 2.2 689 0.8 0.5 1.5 5536 4.3 1.1 5.9 18807 5.1 1.8 9.0 40N_5K_A 515.0 799 0.5 0.5 1.2 1276 2.0 1.5 3.8 1055 1.5 1.3 3.0 8613 6.9 0.6 8.0 25871 8.1 3.1 13.5 40N_5K_B 556.8 1023 0.6 1.4 2.3 1648 2.2 0.9 3.3 1370 1.7 1.3 3.3 8184 7.2 2.2 10.0 27449 8.6 8.9 19.9 40N_5K_C 565.7 626 0.3 0.3 0.9 832 1.2 0.3 1.6 760 1.1 0.2 1.5 6608 5.2 1.0 6.7 22313 6.8 1.9 10.7 44N_4K_A 580.5 1242 0.9 1.1 2.2 2015 2.6 0.8 3.6 1584 2.6 0.9 3.8 16698 12.8 2.6 16.5 65445 19.4 7.0 32.1 44N_4K_B 597.3 940 0.5 0.5 1.3 1700 1.8 0.2 2.3 1312 1.9 0.7 2.8 11465 9.7 1.5 12.3 32793 9.5 5.2 17.6 44N_4K_C 588.9 1596 1.0 4.4 5.8 2653 3.2 2.0 5.4 1978 3.5 2.6 6.4 15657 16.5 3.0 20.8 61526 19.3 11.9 36.7 44N_5K_A 582.1 1162 0.7 0.5 1.6 1771 2.0 0.7 2.9 1460 2.0 0.4 2.6 10167 7.5 0.8 8.9 32297 10.2 1.9 14.9 44N_5K_B 624.9 945 0.6 1.0 1.9 1426 1.8 0.7 2.7 1143 1.8 1.5 3.5 10108 7.6 0.6 8.9 39795 12.8 4.0 20.2 44N_5K_C 592.1 1056 0.6 2.0 2.8 1313 2.4 1.0 3.7 1237 1.7 1.2 3.1 10907 7.5 1.8 10.1 38224 11.8 5.1 20.2 50N_4K_A 665.0 1570 1.1 2.3 3.9 2997 4.7 2.8 7.9 2273 4.7 6.9 12.1 12507 11.7 3.8 16.5 41095 13.4 15.7 34.6 50N_4K_B 685.3 1322 0.7 0.7 1.9 2094 3.2 0.4 3.9 1732 3.3 0.4 4.0 18800 13.8 0.8 15.9 59798 29.0 6.0 41.5 50N_4K_C 641.7 1146 0.8 2.2 3.4 1888 2.8 1.3 4.3 1525 2.8 1.3 4.5 14709 10.8 2.2 14.0 46530 18.4 17.5 41.0 50N_5K_A 645.5 2245 2.3 10.2 13.2 5473 7.2 4.5 12.1 3267 9.1 5.3 14.9 32723 26.7 5.8 34.5 124110 52.0 87.2 152.0 50N_5K_B 666.2 1393 0.8 3.6 5.0 2623 4.5 2.1 7.0 1896 3.9 1.2 5.5 21982 16.3 3.3 21.5 71971 25.8 12.9 46.2 50N_5K_C 667.8 1163 0.8 4.7 5.9 1438 3.1 6.8 10.3 1347 2.6 4.4 7.4 18851 13.5 4.6 19.5 77905 24.8 38.4 72.9 60N_4K_A 814.5 1830 1.2 7.2 8.9 3313 4.8 4.5 9.7 2514 4.8 4.7 9.9 30150 23.0 4.0 29.3 88197 34.2 23.1 66.1 60N_4K_B 727.3 4533 4.5 21.2 26.9 10314 18.3 16.3 35.5 5961 18.0 15.0 33.7 68432 58.2 10.5 72.8 257386 86.3 142.3 252.1 60N_4K_C 747.1 2588 1.7 16.2 18.7 5205 12.6 4.8 18.2 3409 9.2 5.8 15.6 40287 32.3 6.2 41.3 141157 43.3 58.6 117.5 60N_5K_A 743.2 2618 1.7 6.4 9.1 5543 12.9 0.9 14.5 3726 8.8 2.1 11.4 44007 35.9 4.9 43.8 146211 45.4 33.5 96.0 60N_5K_B 800.8 2260 1.7 2.8 5.4 4585 9.2 3.6 13.4 3103 8.1 1.8 10.5 39063 30.2 3.8 37.0 126061 41.1 62.5 115.6 60N_5K_C 822.7 2125 2.0 6.4 9.1 3647 6.7 1.7 8.9 2633 6.6 2.8 9.8 34999 27.5 5.3 35.4 128702 42.4 21.8 75.8 Avg 1301.5 3.4 4.8 2328.4 2.1 6.3 1702.9 2.3 6.2 17471.0 2.6 18.0 60836.1 19.5 46.0 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; CPU: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver. Avg represents the average metric values calculated across all instances.
@@
Table 4: Computational performance of PSFF with different fragment sets for DARP-LPT within 30 minutes (Riedler & Raidl’s instances, $L=4$)
$S_{RF}$ $S_{FF}$ $S_{MF}$ $S_{EFF}$ $S_{ERF}$ Name OBJ F Net CPU Time F Net CPU Time F Net CPU Time F Net CPU Time F Net CPU Time 30N_4K_A 392.2 376 0.2 0.5 1.3 386 1.2 1.1 2.4 411 0.7 0.3 1.2 4188 3.6 0.8 5.0 12876 5.1 1.3 12.0 30N_4K_B 459.7 179 0.1 0.1 0.3 174 0.6 0.4 1.4 180 0.2 0.2 0.4 1775 1.4 0.5 2.1 6404 1.8 0.4 3.1 30N_4K_C 419.5 603 0.3 20.9 21.8 2565 11.4 15.8 27.5 777 5.4 4.9 10.4 13862 18.4 11.2 31.3 17664 8.3 58.8 69.8 30N_5K_A 408.5 525 0.3 4.3 4.9 1396 2.7 10.9 13.8 694 2.3 5.9 8.5 9187 9.1 2.5 12.9 13595 4.8 17.0 24.8 30N_5K_B 425.2 327 0.3 1.2 1.7 499 1.0 1.4 2.6 366 1.1 3.6 5.0 4269 7.5 1.9 10.8 8955 2.6 4.4 9.3 30N_5K_C 445.1 466 0.3 1.7 2.3 857 1.6 1.7 3.5 564 2.2 1.7 4.1 5210 9.4 1.8 12.3 11851 4.5 11.2 17.5 40N_4K_A 525.3 759 0.5 3.8 4.8 2277 7.0 3.7 11.1 918 5.8 2.4 8.5 15364 25.8 7.9 36.0 27638 8.8 80.9 95.1 40N_4K_B 522.2 998 0.7 3.8 5.0 1752 8.7 3.1 12.2 1288 6.8 3.3 10.5 25987 29.0 10.3 42.4 52456 18.0 26.4 52.7 40N_4K_C 557.5 629 0.5 7.5 8.6 822 2.1 4.2 6.5 696 1.9 9.0 11.2 6472 7.2 1.6 10.1 18807 6.1 19.9 30.1 40N_5K_A 477.5 799 0.5 1.6 2.7 3252 10.0 1.8 12.3 1069 12.4 2.3 15.0 20806 27.0 2.2 31.8 25871 12.7 19.6 38.1 40N_5K_B 528.9 1023 0.7 12.0 13.2 3705 7.6 23.3 31.5 1397 7.7 17.6 25.7 16061 19.9 13.6 35.8 27449 8.9 79.8 94.7 40N_5K_C 536.8 626 0.3 0.4 1.1 1768 4.0 0.3 4.7 769 3.9 0.4 4.6 12393 14.8 0.8 17.4 22313 6.7 10.6 21.6 44N_4K_A 541.6 1242 0.9 4.5 5.9 6294 14.4 4.2 19.4 1623 13.8 4.3 18.5 50962 54.8 10.6 70.6 65445 29.0 185.5 225.2 44N_4K_B 562.4 940 0.6 8.1 9.4 6362 9.6 10.2 20.5 1341 13.2 11.1 24.7 32137 33.7 6.1 43.0 32793 12.3 155.9 174.3 44N_4K_C 551.6 1596 1.1 32.2 33.9 12758 29.5 19.6 50.3 2125 33.3 25.5 59.4 61590 78.3 43.1 127.7 61526 23.0 247.0 279.7 44N_5K_A 554.9 1162 0.7 13.3 14.7 4183 7.2 15.7 23.4 1491 8.6 8.6 17.7 23984 25.7 5.0 33.4 32297 13.0 142.7 161.3 44N_5K_B 589.3 945 0.6 10.7 11.9 3677 8.7 9.8 19.3 1167 8.4 9.9 18.7 22135 28.0 4.4 34.7 39795 15.4 19.2 40.8 44N_5K_C 568.4 1056 0.7 26.6 27.9 2218 6.6 22.7 29.8 1277 6.3 24.7 31.4 23095 25.7 16.4 44.7 38224 12.3 105.6 125.6 50N_4K_A 628.4 1570 1.3 31.1 33.3 7105 18.3 45.2 64.2 2310 16.3 28.3 45.1 24443 39.2 12.1 54.2 41095 15.3 355.3 378.9 50N_4K_B 639.0 1322 0.8 6.0 7.6 7106 20.3 5.0 26.2 1755 18.2 4.0 22.7 51665 53.1 7.5 65.8 59798 21.5 172.1 203.3 50N_4K_C 606.8 1146 0.7 13.7 15.1 5169 13.5 15.3 29.6 1530 11.3 14.6 26.4 30064 32.4 15.7 51.6 46530 15.3 500.0 524.0 50N_5K_A 601.8 2245 1.7 177.2 179.9 45134 120.3 294.2 418.0 3506 108.9 127.1 236.8 221956 322.3 211.4 553.3 124110 53.7 1723.0 1800.9 50N_5K_B 624.4 1393 1.0 55.3 57.0 11381 29.3 32.0 62.9 1935 29.0 26.7 56.2 81754 99.2 102.5 212.1 71971 26.2 399.7 439.6 50N_5K_C 619.7 1163 0.7 46.2 47.6 2141 11.6 20.5 32.5 1362 11.1 22.4 34.4 33933 35.0 32.6 71.1 77905 32.9 289.3 338.5 60N_4K_A 762.8 1830 1.6 16.3 19.7 10331 35.1 15.1 51.3 2592 33.0 5.9 39.6 93547 133.0 57.8 200.7 88197 29.0 315.9 364.2 60N_4K_B 680.1 4533 5.9 205.6 214.8 103503 273.4 337.3 619.1 6569 273.1 113.5 388.1 553669 797.0 586.0 1432.7 No 60N_4K_C 698.1 2588 2.1 270.2 273.7 32289 91.7 234.5 329.3 3642 88.7 198.2 288.0 212470 303.8 147.1 472.8 No 60N_5K_A 698.8 2618 2.3 37.1 41.2 36234 107.6 78.8 189.2 4011 105.0 27.2 133.1 229204 327.9 244.7 594.6 No 60N_5K_B 748.1 2260 1.7 133.2 136.6 25308 65.7 152.4 220.3 3234 63.8 128.8 193.4 166511 245.4 203.4 467.3 No 60N_5K_C 768.6 2125 1.6 167.4 170.7 14714 51.2 107.4 160.1 2878 52.3 96.6 149.6 151402 222.1 169.3 407.0 128702 44.6 505.3 572.4 Avg 1301.5 43.8 45.6 11845.3 49.6 83.2 1782.6 31.0 63.0 73336.5 64.4 172.8 26/30 solved 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; CPU: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver. Avg represents the average metric values calculated across all solvable instances. “No” indicates that no optimal solution is obtained within the 30-minute time limit.
Table 5: Computational performance of PSFF with different fragment sets for DARP-LPT within 30 minutes (Riedler & Raidl’s instances, $L=6$)
$S_{RF}$ $S_{FF}$ $S_{MF}$ Name OBJ F Net CPU Time F Net CPU Time F Net CPU Time 30N_4K_A 267.1 146 10.7 1.4 13.3 134 8.1 0.9 9.5 156 8.1 0.7 9.2 30N_4K_B 313.1 72 8.9 1.0 10.4 66 7.2 0.4 7.8 72 7.2 0.4 9.9 30N_4K_C 280.4 146 9.0 3.3 12.8 132 7.3 4.0 12.1 149 7.0 2.1 9.4 30N_5K_A 251.1 194 11.8 2.2 15.7 205 12.4 1.2 14.1 219 12.0 1.6 14.1 30N_5K_B 295.5 150 10.4 1.3 12.2 140 8.2 0.5 9.2 155 7.2 0.8 8.5 30N_5K_C 299.4 180 8.4 1.2 10.3 195 7.9 1.6 9.9 206 7.2 4.8 12.4 40N_4K_A 371.2 259 12.0 2.5 16.0 266 7.6 1.7 10.1 273 11.0 1.5 13.6 40N_4K_B 311.3 441 10.6 9.1 21.7 445 12.0 2.7 15.7 479 17.8 5.2 24.1 40N_4K_C 374.1 197 10.6 0.9 12.8 182 14.3 0.7 15.7 192 8.5 0.9 10.2 40N_5K_A 294.6 271 9.1 2.7 13.1 327 10.0 2.0 13.1 321 8.3 1.5 10.8 40N_5K_B 340.8 327 11.6 2.6 15.8 395 16.8 2.2 19.9 373 7.9 1.9 10.7 40N_5K_C 366.7 229 9.3 6.1 16.6 226 9.5 4.4 14.8 239 7.6 2.5 10.9 44N_4K_A 369.2 361 13.7 4.8 20.4 437 9.6 3.5 14.3 423 11.3 4.9 17.5 44N_4K_B 381.4 383 10.2 10.6 23.0 547 13.2 3.1 17.5 498 15.0 7.0 23.4 44N_4K_C 390.5 503 11.5 20.3 34.0 668 11.6 11.5 24.6 601 12.9 21.3 36.4 44N_5K_A 340.6 533 13.2 4.9 19.9 719 13.1 1.7 16.3 614 12.8 2.7 17.5 44N_5K_B 368.0 326 9.1 6.9 17.4 378 10.5 6.2 18.1 375 10.8 10.4 22.9 44N_5K_C 369.3 334 14.6 9.7 25.6 319 16.0 7.3 24.5 352 11.8 8.1 21.6 50N_4K_A 396.8 475 10.3 13.2 26.0 535 13.5 6.3 21.6 542 23.0 10.9 36.2 50N_4K_B 423.9 539 12.8 11.7 27.5 686 17.9 6.9 26.8 632 21.8 13.3 37.6 50N_4K_C 386.5 519 12.7 9.2 24.3 802 13.5 7.9 23.7 681 19.0 15.4 37.5 50N_5K_A 388.7 962 18.7 28.1 50.6 1928 22.7 36.7 64.4 1283 24.2 31.8 59.9 50N_5K_B 407.2 612 17.5 25.3 45.7 965 17.4 28.7 49.3 769 21.5 48.8 73.2 50N_5K_C 401.2 415 16.1 16.2 34.2 458 12.8 6.1 21.0 473 13.1 9.6 24.7 60N_4K_A 509.6 622 11.2 12.5 27.4 741 18.1 7.6 28.9 728 22.2 8.8 34.7 60N_4K_B 437.0 1502 18.3 88.7 114.9 2657 38.9 58.1 104.5 1879 35.1 26.8 68.5 60N_4K_C 475.4 891 12.9 116.9 135.6 1432 28.1 161.1 194.2 1162 22.8 40.1 67.8 60N_5K_A 476.8 736 10.1 64.9 80.7 1203 21.9 32.5 58.4 985 27.1 92.6 124.5 60N_5K_B 469.5 781 12.5 25.6 43.0 1194 42.8 15.7 63.3 987 29.1 34.6 70.0 60N_5K_C 531.9 535 10.3 12.0 26.3 634 16.4 14.2 33.9 593 20.4 18.2 41.8 Avg 454.7 17.2 31.6 633.9 14.6 31.9 547.0 14.3 32.0 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; CPU: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver. Avg represents the average metric values calculated across all instances.
Table 6: Computational performance of PSFF with different fragment sets for MDARP-LPT within 30 minutes (Riedler & Raidl’s instances, $L=4$)
Appendix C Full computational results of FFF for DARP-LPT
Table 1 describe the computational performance of FFF with different fragment sets for DARP-LPT within 30 minutes.
$S_{RF}$ $S_{FF}$ Name F Net CPU Time OBJ LB Gap F Net CPU Time OBJ LB Gap a3-18 73 0.3 0.0 0.4 302.5 302.5 0.00% 64 0.1 0.0 0.2 302.5 302.5 0.00% a3-30 93 0.1 1799.9 1800.0 494.8 491.0 0.75% 89 0.1 1799.9 1800.0 494.8 491.1 0.74% a3-36 121 0.1 1800.5 1800.0 574.4 560.6 2.40% 112 0.2 1797.6 1800.0 574.4 565.0 1.63% a4-48 318 0.2 1798.9 1800.0 697.9 686.1 1.69% 312 0.5 1799.4 1800.0 697.9 687.4 1.50% a6-60 601 0.4 1793.4 1800.0 869.0 843.2 2.97% 648 1.4 1795.5 1800.0 864.9 843.4 2.48% a7-70 1036 0.8 1796.3 1800.0 986.3 958.5 2.82% 1273 2.6 1792.3 1800.0 986.6 956.5 3.05% a8-80 1088 0.8 1789.1 1800.0 1064.6 1016.4 4.53% 1473 5.1 1785.7 1800.0 1071.1 1019.3 4.84% a8-96 1465 1.0 1756.7 1800.0 1354.8 1296.8 4.28% 1684 7.3 1759.3 1800.0 1353.6 1302.7 3.76% Avg-A 599.4 1566.9 1575.1 2.43% 706.9 1566.2 1575.0 2.25% b3-36 169 0.1 1799.9 1800.0 576.5 566.2 1.79% 161 0.1 1799.7 1800.0 576.5 567.7 1.53% b4-40 311 0.2 1799.5 1800.0 635.2 630.4 0.76% 203 0.4 1799.1 1800.0 635.2 631.1 0.64% b5-50 518 0.4 1798.6 1800.0 789.1 766.2 2.91% 803 1.2 1797.6 1800.0 789.1 768.3 2.63% b6-60 940 0.6 1793.3 1800.0 931.4 895.5 3.86% 1218 2.6 1792.3 1800.0 931.7 900.7 3.33% b7-70 960 0.7 1796.4 1800.0 948.0 921.6 2.79% 1655 3.2 1789.5 1800.0 950.4 920.7 3.13% b8-80 1195 2.0 1790.3 1800.0 1105.8 1071.8 3.08% 1813 5.1 1788.6 1800.0 1104.8 1072.0 2.97% b8-96 No 6540 33.2 1734.9 1800.0 1322.6 1260.9 4.67% Avg-B 682.2 1796.3 1800.0 2.53% 975.5 1794.5 1800.0 2.37% 1 F: fragment count; Net: network time, the time required for fragment generation and fragment-based network construction; MIP: CPU time for solver; Time (s) represents the total time, including preprocessing time, network time, and CPU time for solver. The rows labeled Avg-A and Avg-B represent the average metric values calculated across all instances of type A and type B, respectively. Avg-B does not involve the information of “b8-96” for both $S_{RF}$ and $S_{FF}$ since the former cannot lead to any feasible solution within the 30 minutes.
Table 1: Computational performance of FFF with different fragment sets for DARP-LPT within 30 minutes (Cordeau’s instances, $L=4$)